# Latent Factor Analysis in Short Panels

### v2

@@ E.3 Asymptotic expansions of estimators $\hat{V}_{\varepsilon}$ and $\hat{F}$
and $\tilde{A}=\frac{1}{2}(A-A^{\prime})$ is an antisymmetric $k\times k$ random matrix. To find the antisymmetric matrix $\tilde{A}=(\tilde{a}_{\ell,j})$, we use that $\hat{F}^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}$ is diagonal. Plugging the expansions of the FA estimates, for the term at order $1/\sqrt{n}$ we get that the out-of-diagonal elements of matrix $\Psi_{F}^{\prime}V_{\varepsilon}^{-1}F+F^{\prime}V_{\varepsilon}^{-1}\Psi_{F}-F^{\prime}V_{\varepsilon}^{-1}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F=\frac{1}{2}\Gamma^{-1}F^{\prime}V_{\varepsilon}^{-1}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F+\frac{1}{2}F^{\prime}V_{\varepsilon}^{-1}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+\Gamma\tilde{A}-\tilde{A}\Gamma-F^{\prime}V_{\varepsilon}^{-1}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F$ are nil. Setting the $(\ell,j)$ element of this matrix equal to 00, we get $\tilde{a}_{\ell,j}=-\tilde{a}_{j,\ell}=\frac{1}{\gamma_{j}-\gamma_{\ell}}\left[\frac{1}{2}(\frac{1}{\gamma_{j}}+\frac{1}{\gamma_{\ell}})F^{\prime}_{\ell}V_{\varepsilon}^{-1}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}-F_{\ell}^{\prime}V_{\varepsilon}^{-1}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F_{j}\right]$, for $j\neq\ell$. Then, from Equation (E.10), the $j$th column of $\Psi_{F}$ is $\Psi_{F_{j}}=\frac{1}{\gamma_{j}}M_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}+\frac{1}{2\gamma_{j}}P_{F_{j},V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}+\sum_{\ell=1:\ell\neq j}^{k}\frac{1}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell},V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}-\sum_{\ell=1:\ell\neq j}^{k}\frac{\gamma_{\ell}}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell},V_{\varepsilon}}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F_{j}$, where we use $P_{F,V_{\varepsilon}}=\sum_{\ell=1}^{k}P_{F_{\ell},V_{\varepsilon}}$. Part (a) follows.
@@
$diag(\Psi_{y})=diag\left(\sum_{j=1}^{k}(F_{j}\Psi_{F_{j}}^{\prime}+\Psi_{F_{j}}F_{j}^{\prime})+\Psi_{\varepsilon}\right).$ (E.11)
From part (a) and the definition of $P_{F_{j},{V}_{\varepsilon}}$ we have $\sum_{j=1}^{k}\Psi_{F_{j}}F_{j}^{\prime}=\frac{1}{2}\sum_{j=1}^{k}P_{F_{j},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}+M_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F,{V}_{\varepsilon}}^{\prime}+\sum_{\ell\neq j}\frac{\gamma_{j}}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}-\sum_{\ell\neq j}^{k}\frac{\gamma_{\ell}\gamma_{j}}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell},{V}_{\varepsilon}}\Psi_{\varepsilon}P_{F_{j},{V}_{\varepsilon}}^{\prime}$ $=:N_{1}+N_{2}+N_{3}+N_{4},$ where $P_{F,{V}_{\varepsilon}}=\sum_{j=1}^{k}P_{F_{j},{V}_{\varepsilon}}=I_{T}-M_{F,{V}_{\varepsilon}}$ and $\sum_{\ell\neq j}$ denotes the double sum over $j,\ell=1,...,k$ such that $\ell\neq j$. Matrix $N_{1}$ is symmetric and it contributes $2N_{1}$ to the RHS of (E.11). Instead, matrix $N_{4}$ is antisymmetric (it can be seen by interchanging indices $j$ and $\ell$ in the summation) and it does not contribute to the RHS of (E.11). For matrix $N_{3}$ we have $N_{3}+N_{3}^{\prime}=\sum_{\ell\neq j}\frac{\gamma_{j}}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}+\sum_{\ell\neq j}\frac{\gamma_{\ell}}{\gamma_{\ell}-\gamma_{j}}P_{F_{\ell},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}=\sum_{\ell\neq j}P_{F_{\ell},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}=\sum_{\ell,j}P_{F_{\ell},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}-\sum_{j}P_{F_{j},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}=P_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F,{V}_{\varepsilon}}^{\prime}-2N_{1},$ where we have interchanged $\ell$ and $j$ in the first equality when writing $N_{3}^{\prime}$. Thus, we get:
$\sum_{j=1}^{k}(F_{j}\Psi_{F_{j}}^{\prime}+\Psi_{F_{j}}F_{j}^{\prime})$ $=$ $M_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F,{V}_{\varepsilon}}^{\prime}+P_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})M_{F,{V}_{\varepsilon}}^{\prime}+P_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F,{V}_{\varepsilon}}^{\prime}$ (E.12)
@@
Then, Equation (E.11) with (E.12) yields Equation (E.4).
E.4 Local analysis of the first-order conditions of FA estimators
Consider the criterion $L(\theta)=-\frac{1}{2}\log|\Sigma(\theta)|-\frac{1}{2}Tr\left(V_{y}\Sigma(\theta)\right)$, where $V_{y}$ is a p.d. matrix in a neighbourhood of $V_{y}^{0}$. In our Assumptions, $\theta_{0}$ is an interior point of $\Theta$. Let $\theta^{*}=(vec(F^{*})^{\prime},diag(V_{\varepsilon}^{*})^{\prime})^{\prime}$ denote the maximizer of $L(\theta)$ subject to $\theta\in\Theta$. According to Anderson (2003), the first-order conditions (FOC) for the maximization of $L(\theta)$ are: (a) $diag(V_{y})=diag(F^{*}(F^{*})^{\prime}+V_{\varepsilon}^{*})$ and (b) $F^{*}$ is the matrix of eigenvectors of $V_{y}(V_{\varepsilon}^{*})^{-1}$ associated to the $k$ largest eigenvalues $1+\gamma_{j}^{*}$ for $j=1,...,k$, normalized such that $(F^{*})^{\prime}(V_{\varepsilon}^{*})^{-1}F^{*}=diag(\gamma_{1}^{*},...,\gamma_{k}^{*})$.
@@ E.4 Local analysis of the first-order conditions of FA estimators
In Lemma 6, condition (a) corresponds to Assumption A.5 and is equivalent to condition (b) that $\boldsymbol{X}$ is full-rank. Condition (c) is used in Theorem 5.9 of Anderson and Rubin (1956) to show local identification. Condition (d) involves the second-order partial derivatives of the population criterion function. While the Hessian matrix $J_{0}$ itself is singular because of the rotational invariance of the model to latent factors, the second-order partial derivatives matrix along parameter directions, which are in the tangent plan to the contraint set, is non-singular. Condition (d) is equivalent to invertibility of the bordered Hessian.
@@
From the arguments in the proof of Proposition 7, Equations (E.11) and (E.12), we get:
$\Psi_{F}^{\epsilon}F_{0}^{\prime}+F_{0}(\Psi_{F}^{\epsilon})^{\prime}=\Psi_{y}^{\epsilon}-\Psi_{V_{\varepsilon}}^{\epsilon}-M_{F_{0},V_{\varepsilon}^{0}}(\Psi_{y}^{\epsilon}-\Psi_{V_{\varepsilon}}^{\epsilon})M_{F_{0},V_{\varepsilon}^{0}}^{\prime},$ (E.13)
$diag(M_{F_{0},V_{\varepsilon}^{0}}(\Psi_{y}^{\epsilon}-\Psi_{V_{\varepsilon}}^{\epsilon})M_{F_{0},V_{\varepsilon}^{0}}^{\prime})=0.\qquad\qquad\qquad$ (E.14)
As in Section E.3, Equation (E.14) yields:
@@
$diag(\Psi_{V_{\varepsilon}}^{\epsilon})=(V_{\varepsilon}^{0})^{2}\left(\boldsymbol{X}^{\prime}\boldsymbol{X}\right)^{-1}\boldsymbol{X}^{\prime}vech\left(G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}\right).$ (E.15)
@@
Now, using Equation (E.13), we get $V_{y}-\Sigma^{*}=\epsilon\left(\Psi_{y}^{\epsilon}-F_{0}(\Psi_{F}^{\epsilon})^{\prime}-\Psi_{F}^{\epsilon}F_{0}^{\prime}-\Psi_{V_{\varepsilon}}^{\epsilon}\right)+O(\epsilon^{2})$ $=\epsilon M_{F_{0},V_{\varepsilon}^{0}}(\Psi_{y}^{\epsilon}-\Psi_{V_{\varepsilon}}^{\epsilon})M_{F_{0},V_{\varepsilon}^{0}}^{\prime}+O(\epsilon^{2})=\epsilon G_{0}\Delta^{*}G_{0}^{\prime}+O(\epsilon^{2}),$ where $\Delta^{*}:=G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}-G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{V_{\varepsilon}}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}$. Using that $vech(G_{0}^{\prime}diag(a)G_{0})=\boldsymbol{X}a$, and Equation (E.15), the vectorized form of matrix $\Delta^{*}$ is: $vech(\Delta^{*})=vech\left(G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}\right)$ $-\boldsymbol{X}(V_{\varepsilon}^{0})^{-2}diag(\Psi_{V_{\varepsilon}}^{\epsilon})=M_{\boldsymbol{X}}vech\left(G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}\right).$ Thus, we have shown that, at first order in $\epsilon$, the difference between $V_{y}=V_{y}^{0}+\epsilon\Psi_{y}^{\epsilon}$ and the FA $k$-factor approximation $\Sigma^{*}$ is $\epsilon G_{0}\Delta^{*}G_{0}^{\prime}$, with $vech(\Delta^{*})=M_{\boldsymbol{X}}vech\left(G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}\right).$ It shows that the small perturbation $\epsilon\Psi_{y}^{\epsilon}$ around $V_{y}^{0}$ keeps the DGP within the $k$-factor specification (at first order) if, and only if, we have that vector $vech\left(G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}\right)$ is spanned by the columns of $\boldsymbol{X}$.
@@
Consider $\Psi_{y}^{\epsilon}=H\xi\xi^{\prime}H^{\prime}$, where $H:=[F_{0}:G_{0}]$ and vector $\xi=(\xi_{F}^{\prime},\xi_{G}^{\prime})^{\prime}$ are partitioned in $k$ and $T-k$ dimensional components, which corresponds to a local alternative with $(k+1)$th factor $H\xi$ and small loading $\epsilon$ in the perturbation $\epsilon\Psi_{y}^{\epsilon}$. Then, we have $G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}=\xi_{G}\xi_{G}^{\prime}$ since $F_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}G_{0}=0$ and $G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}G_{0}=I_{T-k}$. Thus, $vech(\Delta^{*})=M_{\boldsymbol{X}}vech\left(\xi_{G}\xi_{G}^{\prime}\right)$. Hence, it is only the component of $vech\left(\xi_{G}\xi_{G}^{\prime}\right)$ that is orthogonal to the range of $\boldsymbol{X}$, which generates a local deviation from a $k$-factor specification through the multiplication by the projection matrix $M_{\boldsymbol{X}}$. It clarifies the role of the projector in the local power. On the contrary, the component spanned by the columns of $\boldsymbol{X}$ can be “absorbed" in the $k$-factor specification by a redefinition of the factor $F$ and the variance $V_{\varepsilon}$ through $F^{*}$ and $V_{\varepsilon}^{*}$.
E.5 Feasible asymptotic normality of the FA estimators
E.5.1 Asymptotic expansions

### v3

@@ B.3 Asymptotic expansions of estimators $\hat{V}_{\varepsilon}$ and $\hat{F}$
and $\tilde{A}=\frac{1}{2}(A-A^{\prime})$ is an antisymmetric $k\times k$ random matrix. To find the antisymmetric matrix $\tilde{A}=(\tilde{a}_{\ell,j})$, we use that $\hat{F}^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}$ is diagonal. Plugging the expansions of the FA estimates, for the term at order $1/\sqrt{n}$ we get that the out-of-diagonal elements of matrix $\Psi_{F}^{\prime}V_{\varepsilon}^{-1}F+F^{\prime}V_{\varepsilon}^{-1}\Psi_{F}-F^{\prime}V_{\varepsilon}^{-1}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F=\frac{1}{2}\Gamma^{-1}F^{\prime}V_{\varepsilon}^{-1}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F+\frac{1}{2}F^{\prime}V_{\varepsilon}^{-1}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+\Gamma\tilde{A}-\tilde{A}\Gamma-F^{\prime}V_{\varepsilon}^{-1}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F$ are nil. Setting the $(\ell,j)$ element of this matrix equal to 0, we get $\tilde{a}_{\ell,j}=-\tilde{a}_{j,\ell}=\frac{1}{\gamma_{j}-\gamma_{\ell}}\left[\frac{1}{2}(\frac{1}{\gamma_{j}}+\frac{1}{\gamma_{\ell}})F^{\prime}_{\ell}V_{\varepsilon}^{-1}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}-F_{\ell}^{\prime}V_{\varepsilon}^{-1}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F_{j}\right]$, for $j\neq\ell$. Then, from Equation (B.10), the $j$th column of $\Psi_{F}$ is $\Psi_{F_{j}}=\frac{1}{\gamma_{j}}M_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}+\frac{1}{2\gamma_{j}}P_{F_{j},V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}+\sum_{\ell=1:\ell\neq j}^{k}\frac{1}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell},V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}-\sum_{\ell=1:\ell\neq j}^{k}\frac{\gamma_{\ell}}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell},V_{\varepsilon}}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F_{j}$, where we use $P_{F,V_{\varepsilon}}=\sum_{\ell=1}^{k}P_{F_{\ell},V_{\varepsilon}}$. Part (a) follows.
@@
$diag(\Psi_{y})=diag\left(\sum_{j=1}^{k}(F_{j}\Psi_{F_{j}}^{\prime}+\Psi_{F_{j}}F_{j}^{\prime})+\Psi_{\varepsilon}\right).$ (B.11)
From part (a) and the definition of $P_{F_{j},{V}_{\varepsilon}}$ we have $\sum_{j=1}^{k}\Psi_{F_{j}}F_{j}^{\prime}=\frac{1}{2}\sum_{j=1}^{k}P_{F_{j},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}+M_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F,{V}_{\varepsilon}}^{\prime}+\sum_{\ell\neq j}\frac{\gamma_{j}}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}-\sum_{\ell\neq j}^{k}\frac{\gamma_{\ell}\gamma_{j}}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell},{V}_{\varepsilon}}\Psi_{\varepsilon}P_{F_{j},{V}_{\varepsilon}}^{\prime}$ $N_{1}+N_{2}+N_{3}+N_{4},$ where $P_{F,{V}_{\varepsilon}}=\sum_{j=1}^{k}P_{F_{j},{V}_{\varepsilon}}=I_{T}-M_{F,{V}_{\varepsilon}}$ and $\sum_{\ell\neq j}$ denotes the double sum over $j,\ell=1,...,k$ such that $\ell\neq j$. Matrix $N_{1}$ is symmetric and it contributes $2N_{1}$ to the RHS of (B.11). Instead, matrix $N_{4}$ is antisymmetric (it can be seen by interchanging indices $j$ and $\ell$ in the summation) and it does not contribute to the RHS of (B.11). For matrix $N_{3}$ we have $N_{3}+N_{3}^{\prime}=\sum_{\ell\neq j}\frac{\gamma_{j}}{\gamma_{j}-\gamma_{\ell}}P_{F_{\ell},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}+\sum_{\ell\neq j}\frac{\gamma_{\ell}}{\gamma_{\ell}-\gamma_{j}}P_{F_{\ell},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}=\sum_{\ell\neq j}P_{F_{\ell},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}=\sum_{\ell,j}P_{F_{\ell},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}-\sum_{j}P_{F_{j},{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F_{j},{V}_{\varepsilon}}^{\prime}=P_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F,{V}_{\varepsilon}}^{\prime}-2N_{1},$ where we have interchanged $\ell$ and $j$ in the first equality when writing $N_{3}^{\prime}$. Thus, we get:
$\sum_{j=1}^{k}(F_{j}\Psi_{F_{j}}^{\prime}+\Psi_{F_{j}}F_{j}^{\prime})$ $=$ $M_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F,{V}_{\varepsilon}}^{\prime}+P_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})M_{F,{V}_{\varepsilon}}^{\prime}+P_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})P_{F,{V}_{\varepsilon}}^{\prime}$ (B.12)
@@
Then, Equation (B.11) with (B.12) yields Equation (B.4).
B.4 Local analysis of the first-order conditions of FA estimators
Consider the criterion $L(\theta)=-\frac{1}{2}\log|\Sigma(\theta)|-\frac{1}{2}Tr\left(V_{y}\Sigma(\theta)\right)$, where $V_{y}$ is a p.d. matrix in a neighbourhood of $V_{y}^{0}$. In our Assumptions, $\theta_{0}$ is an interior point of $\Theta$. Let $\theta^{*}=(vec(F^{*})^{\prime},diag(V_{\varepsilon}^{*})^{\prime})^{\prime}$ denote the maximizer of $L(\theta)$ subject to $\theta\in\Theta$. According to Anderson (2003), the first-order conditions (FOC) for the maximization of $L(\theta)$ are: (a) $diag(V_{y})=diag(F^{*}(F^{*})^{\prime}+V_{\varepsilon}^{*})$ and (b) $F^{*}$ is the matrix of eigenvectors of $V_{y}(V_{\varepsilon}^{*})^{-1}$ associated to the $k$ largest eigenvalues $1+\gamma_{j}^{*}$ for j=1,,kj=1,...,k, normalized such that $(F^{*})^{\prime}(V_{\varepsilon}^{*})^{-1}F^{*}=diag(\gamma_{1}^{*},...,\gamma_{k}^{*})$.
@@ B.4 Local analysis of the first-order conditions of FA estimators
In Lemma 7, condition (a) corresponds to Assumption 7 and is equivalent to condition (b) that $\boldsymbol{X}$ is full-rank. Condition (c) is used in Theorem 5.9 of Anderson and Rubin (1956) to show local identification. Condition (d) involves the second-order partial derivatives of the population criterion function. While the Hessian matrix $J_{0}$ itself is singular because of the rotational invariance of the model to latent factors, the second-order partial derivatives matrix along parameter directions, which are in the tangent plan to the contraint set, is non-singular. Condition (d) is equivalent to invertibility of the bordered Hessian.
@@
From the arguments in the proof of Proposition 6, Equations (B.11) and (B.12), we get:
$\Psi_{F}^{\epsilon}F_{0}^{\prime}+F_{0}(\Psi_{F}^{\epsilon})^{\prime}=\Psi_{y}^{\epsilon}-\Psi_{V_{\varepsilon}}^{\epsilon}-M_{F_{0},V_{\varepsilon}^{0}}(\Psi_{y}^{\epsilon}-\Psi_{V_{\varepsilon}}^{\epsilon})M_{F_{0},V_{\varepsilon}^{0}}^{\prime},$ (B.13)
$diag(M_{F_{0},V_{\varepsilon}^{0}}(\Psi_{y}^{\epsilon}-\Psi_{V_{\varepsilon}}^{\epsilon})M_{F_{0},V_{\varepsilon}^{0}}^{\prime})=0.\qquad\qquad\qquad$ (B.14)
As in Section B.3, Equation (B.14) yields:
@@
$diag(\Psi_{V_{\varepsilon}}^{\epsilon})=(V_{\varepsilon}^{0})^{2}\left(\boldsymbol{X}^{\prime}\boldsymbol{X}\right)^{-1}\boldsymbol{X}^{\prime}vech\left(G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}\right).$ (B.15)
@@
Now, using Equation (B.13), we get $V_{y}-\Sigma^{*}=\epsilon\left(\Psi_{y}^{\epsilon}-F_{0}(\Psi_{F}^{\epsilon})^{\prime}-\Psi_{F}^{\epsilon}F_{0}^{\prime}-\Psi_{V_{\varepsilon}}^{\epsilon}\right)+O(\epsilon^{2})$ $\epsilon M_{F_{0},V_{\varepsilon}^{0}}(\Psi_{y}^{\epsilon}-\Psi_{V_{\varepsilon}}^{\epsilon})M_{F_{0},V_{\varepsilon}^{0}}^{\prime}+O(\epsilon^{2})=\epsilon G_{0}\Delta^{*}G_{0}^{\prime}+O(\epsilon^{2}),$ where $\Delta^{*}:=G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}-G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{V_{\varepsilon}}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}$. Using that $vech(G_{0}^{\prime}diag(a)G_{0})=\boldsymbol{X}a$, and Equation (B.15), the vectorized form of matrix $\Delta^{*}$ is: $vech(\Delta^{*})=vech\left(G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}\right)$ $\boldsymbol{X}(V_{\varepsilon}^{0})^{-2}diag(\Psi_{V_{\varepsilon}}^{\epsilon})=M_{\boldsymbol{X}}vech\left(G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}\right).$ Thus, we have shown that, at first order in $\epsilon$, the difference between $V_{y}=V_{y}^{0}+\epsilon\Psi_{y}^{\epsilon}$ and the FA $k$-factor approximation $\Sigma^{*}$ is $\epsilon G_{0}\Delta^{*}G_{0}^{\prime}$, with $vech(\Delta^{*})=M_{\boldsymbol{X}}vech\left(G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}\right).$ It shows that the small perturbation $\epsilon\Psi_{y}^{\epsilon}$ around $V_{y}^{0}$ keeps the DGP within the $k$-factor specification (at first order) if, and only if, we have that vector $vech\left(G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}\right)$ is spanned by the columns of $\boldsymbol{X}$.
@@
Consider $\Psi_{y}^{\epsilon}=H\xi\xi^{\prime}H^{\prime}$, where $H:=[F_{0}:G_{0}]$ and vector $\xi=(\xi_{F}^{\prime},\xi_{G}^{\prime})^{\prime}$ are partitioned in $k$ and TkT-k dimensional components, which corresponds to a local alternative with $(k+1)$th factor $H\xi$ and small loading $\epsilon$ in the perturbation $\epsilon\Psi_{y}^{\epsilon}$. Then, we have $G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}\Psi_{y}^{\epsilon}(V_{\varepsilon}^{0})^{-1}G_{0}=\xi_{G}\xi_{G}^{\prime}$ since $F_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}G_{0}=0$ and $G_{0}^{\prime}(V_{\varepsilon}^{0})^{-1}G_{0}=I_{T-k}$. Thus, $vech(\Delta^{*})=M_{\boldsymbol{X}}vech\left(\xi_{G}\xi_{G}^{\prime}\right)$. Hence, it is only the component of $vech\left(\xi_{G}\xi_{G}^{\prime}\right)$ that is orthogonal to the range of $\boldsymbol{X}$, which generates a local deviation from a $k$-factor specification through the multiplication by the projection matrix $M_{\boldsymbol{X}}$. It clarifies the role of the projector in the local power. On the contrary, the component spanned by the columns of $\boldsymbol{X}$ can be “absorbed" in the $k$-factor specification by a redefinition of the factor $F$ and the variance $V_{\varepsilon}$ through $F^{*}$ and $V_{\varepsilon}^{*}$.
B.5 Feasible asymptotic normality of the FA estimators
B.5.1 Asymptotic expansions